OpenStudy (anonymous):

Suppose that y = 21(1.07)x, represents book sales in billions of dollars since 2000. X represents the number of years after 2000. Predict the book sales for 2010 (x = 10), to the nearest billion. $41 billion $39 billion $44 billion $36 billion

OpenStudy (texaschic101):

y = 21(1.07)^10 y = 21(1.97) y = 41.37 closest to 41 billion